OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)?

OpenStudy (anonymous):

Basically, Put it into slope-intercept form. y = mx + b where m is the slope and b is the y-intercept. First find the slope. 5 + 1/4 + 4 = 6/8 = 3/4 Plug it in for m. y = 3/4x + b Now plug in one of the points in for x and y to solve for b. -1 = 3/4(-4) + b -1 = -3 + b b = 2 y = 3/4x + 2 Since 2, is the y-intercept, 2 or (0, 2) is the answer. :)

OpenStudy (anonymous):

Thanks soo much! I get really confused on these types of questions!

OpenStudy (anonymous):

You're welcome ^_^